The following data set represents the number of NBA games in January 2016, watched by 10 randomly selected student in STAT 206. 7, 0, 4, 2, 2, 1, 0, 1, 2, 3 (a) What is the sample mean? (b) Calculate sample variance. (c) Estimate the mean number of NBA games watched by a student in January 2016. (d) Estimate the standard error of the estimated mean.
Answer #1

(a)

Sample mean = ar{x} = 22/10 = 2.2

(b)

x x - ar{x} (x - ar{x} )2
7 4.8 23.04
0 - 2.2 4.84
4 1.8 3.24
2 - 0.2 0.04
2 - 0.2 0.04
1 - 1.2 1.44
0 - 2.2 4.84
1 - 1.2 1.44
2 - 0.2 0.04
3 0.8 0.64
Total = 39.6

Sample variance = 39.6/9 = 4.4

(c)

Mean number = 2.2

(d)

Standard deviation = s = V4.4- 2.0976

Standard error = s/sqrt{n}

= 2.0976/V10 = 0.6633